# The unified cross-disciplinary model of the operation of neurons

### v2

@@ 4.6 Heat emission/absorption
In the framework of the classic theory, Ohmic currents flow through resistors that dissipate heat due to friction, no matter in which direction ($W=I^{2}*R$) the ion currents flow. The existence of "cooling", alone, undermines the credibility of the classic theory [38]. The issue here is that the dissipation was attributed to the "leakage current." As Hodgkin wrote: "Hill and his colleagues found [71] that an initial phase of heat liberation [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." [72] "All authors came to similar conclusions: during the action potential, no significant net heat is produced. Transient heat releases are mostly reabsorbed in the second phase of the action potential. …The finding of a reversible heat release during the action potential of nerves is a striking and very fundamental fact. It is inconsistent with the HH model. The physics underlying the nervous impulse must rather be based on reversible processes". [73] Our Speculations in Appendix A on the energy consumption of the nervous impulse underpins that the complex physical process comprises mostly reversible disciplinary processes: most of the energy of AP is stored as reversible elastic potential energy. (Alternatively, it may be hypothesized that the positive $K^{+}$ ions, while traveling ahead of the positive $Na^{+}$ ions under the effect of a potential-free space in the second phase of AP, not only turn with their positive charge the offset voltage to negative, but on the way the $K^{+}$ ions collect the heat dissipated by the $Na^{+}$ ions. "Laws of science cannot describe life".)
The heat liberation and absorption are direct experimental evidence of the "slow current" and the mechanisms described above. The lack of leakage current (see section 2.6.4) not only validates our model, but also solves the long-standing mystery of reversible heat release during the AP. Pressure changes accompany the electric process and naturally explain, quantitatively, the observed temperature changes. As our explanation shows, only a small fragment of energy is stored in the form of electric potential energy. Hence, its dissipation is also negligible: "no significant net heat is produced" [73]. The energy is almost exclusively stored as reversible elastic energy. The conclusion in [35], that "brain cells communicate with mechanical pulses, not electric signals", although shocking for the first reading, and is only half of the truth, is essentially correct. More precisely, mechanical pulses (pressure waves) deliver the charge carriers of the electric signals, in close interaction with the repulsion of the delivered charged carriers. In Equation (7), we also must consider the mechanical force delivered by the pressure wave. Furthermore, the nervous pulse transmission over the axons must be revisited. The enormous pressure, the incompressible electric fluid, and the repulsion of the delivered particles suggest an alternative and more reasonable mechanism.
@@ 4.7 Neural entropy
The theoretical diagram line in Fig. 14, based on the presented correct physical model, might serve as a good starting point for interpreting entropy for neuronal operation, again in analogy with the Carnot cycle; although the charge of the ions (the macroscopic current), complicates the process. Furthermore, material convection happens in different ill-defined volume parts of the membrane in different directions. Initially, in the resting state, the ions are in a (relatively) disordered state. After the rush-in, in the dynamic layer, the ions have a highly ordered macroscopic speed (see Stokes-Einstein speed, the gradually changing gradients, and the finite speed) component toward the AIS. In the layer, the maximum ordered state is reached when the voltage drop on the AIS reaches its peak value (the current is slow). Then, the ions slow down, and at zero voltage gradient, they reach again a minimally ordered state. Then, due to the ‘capacitive current’, the current changes its direction, and the backward current represents an ordered state again. Perhaps the latter-mentioned negative contribution is the one referred to as ‘negative entropy’ when describing life through thermodynamic processes? We must be aware that in the background, energy-producing processes are going on, comprising ordered material transport. Considering the neuron, based only on the described primary processes, would be a mistake.
@@
The above discussion might help in understanding the physical entropy in the neuron. However, the information entropy is different. As our discussion underpins, the "information" arrives at the neuronal inputs in the form of charge carriers, and the time course of the input currents, in cooperation with the neural environment as well as the neuron itself, carries the information in some form. The information is distributed in space and time [74, 75], so we can be sure that using the pulse count as a measure of neural information is wrong. We must distinguish the carrier of the information, the ion, and the information content that its temporal appearance carries. After having the correct physical model of operation, we have a chance to gain more knowledge about how the information is encoded in the spatially and temporally distributed observable signals; furthermore, we have better hopes to understand the neural information processing, including the brain’s entropy handling.
5 Controling neuronal stages
From control theory, it is known that the goal of a controlled system is to govern the application of system inputs to drive the system to a desired state while minimizing delay, overshoot, steady-state error, and ensuring control stability. The neuron implements a controller that monitors the controlled process variable (membrane voltage) and compares it with the reference or setpoint (the resting potential). Recall that the geometry, and the composition and concentration of electrolytes, define the setpoint. The difference between the process variable’s actual and desired values, known as the error signal, represents the actual offset potential. It is applied as feedback to generate a control action to bring the controlled process variable to the same value as the setpoint.
5.1 Neuron as a PID controller
Biology uses such a simple Proportional-Integral-Derivative (PID) controller which has a proportional, an integrated and a derivative component. In simple words, it means that the output of the system (the output voltage measurable on the AIS) depends on the input voltage (input current if the resistance is constant), the integrated input current(s) and their derivatives (provided that the membrane’s capacity is constant). The sum of those summands defines the resulting output voltage. The biological case is more complicated than the technical one because biology works with slow ions and so those components have their temporal dependence, furthermore, the summands have several constituents and the neurons have an autonomous processing of those current-related constituents. That is, unlike assumed in the classic physiology, the output voltage is not necessarily a simple function of the mentioned input variables. The classic models all omit the derivatives’ contribution (a direct consequence of using ’clamping’) except that of the capacitive current. Furthermore, the ’Integrate and Fire’-type models consider the integrated currents only in the ’chargeup’ (’Computing’) period of operation and they entirely omit that different physical processes are going on in different stages of operation. That is, even the best classic models can not describe neuronal operation completely and correctly.
The gradients are used to adjust the process variable by their positive and negative contributions (corresponding to rising and falling currents), and the different speeds of the thermodynamic and electrical interactions minimize the delay (i.e., provide the maximum speed of operations, vital for survival). The steady-state error is minimized by setting the process variable to the reference point using long-term stable parameters (geometry and overall concentration). The low-intensity current through the always-open resting ion channels provides dynamic stability in the steady state.
We considered that the neuron has a stable base state. On the one side, this resting state must be dynamically stabilized for little perturbations using as little energy as possible (and to provide a mechanism when the cell grows, divides, or ages). On the other side, it must be able to restore the state after rough perturbations as quickly as possible, causing short-time transients (when restoring the membrane’s potential after issuing a spike). Both in the resting state and in the transient state, the system attempts to return to its balanced state, although this action in the two states require different physical mechanisms.
The system aims to provide an intense output that informs the downstream neurons. This overshoot is implemented by a large number of voltage-gated ion channels (distinct from the non-gated ion channels used to maintain the resting potential) distributed in the neuron’s membrane. The overshoot current flows through those persistently open ion channels, concentrated in AIS (which has about two orders of magnitude higher channel density than the wall). The intense slow current produces a condenser-like behavior (capacitive current); the phenomena called "polarization" and "hyperpolarization" (not polarization; instead, a movement of completely separated charge in the surface layer of the electrolyte) of the membrane provide the necessary positive and negative error signals for the controller in the transient state by moving the actual potential value of the membrane above and below the resting potential value. The potential’s and the electric field’s magnitudes depend on the concentration and the geometry (the finite size of the membrane). The ions’ chemical nature comes into play only if the polarizability can differ for different molecules.
The overall control function of the general PID controller in Fig. 16 is described by
where $T_{i}=\frac{K_{p}}{K_{i}}$ is the time integration constant, $T_{d}=\frac{K_{d}}{K_{p}}$ is the derivative time constant. That is, in the world of a neuron, the "theory of everything" is
$V_{M}^{OUT}(t)=K_{p}\biggl(\overbrace{\underbrace{\overbrace{V_{M}^{IN}(t)}^{external}}_{clamping\ etc}+\overbrace{\underbrace{\frac{1}{T_{i}}\int_{0}^{t}V_{Wall}^{IN}(\tau)d\tau}_{membrane^{\prime}s\ wall}}^{integral,\ parallel\ RC}}^{Hodgkin-Huxley}+\overbrace{\underbrace{T_{d}\frac{dV_{M}^{IN}(t)}{dt}}}^{differential,\ serial\ RC}_{AIS}\,\biggr)$ (47)
Compare it to Eq. (40), which describes the serial $RC$ oscillator, and the usual equations describing HH’s hypothesis. The presence of some terms depends on the neuron’s actual stage; furthermore, the $R$ values in the parallel and serial $RC$ circuit differs. As discussed in section 4.1, subthreshold excitations provide direct experimental evidence that the description of the resting state includes both serial and parallel contributions.
The first term represents the constant effect of the external world (including synaptic inputs, clamping, and exciting brain tissue). The second term represents the resting state when no gradients are present. The voltage is mainly due to current through ion channels in the membrane’s wall; it represents a parallel $RC$ circuit. The third term represents the sum of the gradients due to all effects (the external world, the internal processes, and the outflow through the AIS); it represents a serial $RC$ circuit. HH considered only the first two terms plus the voltage gradient due to the condenser (but not the capacitive current). They missed that the input currents also produce a voltage gradient (clamping, per definitionem, supresses voltage gradients of the membrane) and that the output current through the AIS also produces a gradient. Furthermore, the equation explains that the external currents (say, step-functions’ rising and falling edges) also generate enormous changes in the output voltage. Similarly, an apparently "foreign" invasion (changing mechanically the location of charges by pressure, changing the concentration in one of the segments, ultrasound, magnetic pulse, and so on) generates a change in the concentration or position of charges on the neuronal membrane. This way, it generates a voltage gradient (as well as pressure gradient, furthermore, other changes), and so it may generate neuronal spikes.
From a control theoretical point of view, the condenser geometry and the ion concentrations set the ’always the same’ value (the set-point; see also Figure 1) and the currents serve to keep or restore that value. After introducing a finite-width membrane into an electrolyte, a potential difference between the two membrane surfaces is created; see Eq. (19). When adjusting the membrane’s potential, we must consider its ground and excited states separately, given that the perturbations in these two cases are vastly different. Nature employs various mechanisms to maintain the ground state and recover it after generating a spike. Although nature’s tools are remarkably similar (channels and pumps; here, for discussing ’downhill’ channels, we must consider the charge-up of different components), the significant difference in current intensity needs a different discussion.
The classical condenser has an electric field, as the blue dashed line shows in Fig. 4. No field is outside the condenser or inside the condenser plates (see Fig. 3). There is a sudden jump in the field on the conductor’s surface (the armatures) and a constant field between the plates. (The charged layer can be ideally thin if electrons are sitting on the surface, but is of finite thickness on the opposite side where ions are on the surface. This latter effect is not discussed in classical electricity.)
The biological condenser behaves differently; see the red diagram line. An electrolyte (instead of an insulator) outside the condenser significantly alters the field’s structure. The electric field on the surface and between the plates changes by a factor of 3 to 4. Outside the plates, the polarization creates a field that changes logarithmically. An ion layer with finite thickness is built near the membrane’s surface. It is uniformly charged, so the field inside it is linear. At the internal surface of the condenser, it takes the value of the field calculated for the gap; on the other side of the layer, it takes the value of the logarithmic field at that position. The dashed line represents the gap field, the continuous line represents the charged layer’s field, and the dotted line represents the electrolyte’s field. We can observe the "correspondence principle": the discrete-continuous fields join seamlessly; furthermore, if the polarizability of the electrolyte decreases toward zero, the field value approaches the classical value.
The thickness of the charged layer has a significant impact on the resulting electric field (i.e., the neuron’s ’setpoint’). Suppose biological fragments are formed and settle on the membrane’s intracellular side due to the increased $\Delta z$. In that case, the resting potential (and the threshold potential) may change, potentially leading to neurological diseases. As [76] observed, the setpoint changes with temperature (the "thermodynamic electrical field" changes with the temperature, while the "electrical electrical field" does not, so their crossing point changes), reflected in the dynamics of the AP.
5.2 Robustness of the control circle
An exciting thought experiment is discussed in [5]. "First, we consider a hypothetical membrane that is permeable only to $Na^{+}$ ions. Suppose that $Na^{+}]_{o}$, the outside or extracellular sodium concentration, is 145mM145\ mM, and $Na^{+}]_{i}$ is 12 mM. And suppose that initially, there is no membrane potential. The diffusion gradient for $Na^{+}$ favors $Na^{+}$ entry into the cell, and the initial $Na^{+}$ influx carries a charge that builds up on the inside of the cell. This produces a potential (recall that separation of charge produces a potential) across the membrane that impedes further $Na^{+}$ ion movement because the positive charges repel the positively charged $Na^{+}$ ion."
The paper correctly explains that in the case of $Na^{+}$ only, the potential is set up to 66.6mV66.6\ mV. (We quietly add that plus a layer of the corresponding $Cl^{-}$ ions on the other segment, in a similar concentration ratio, builds up; so also, the attraction between positive and negative charges contributes to the effect.) In the absence of our argumentation, it cannot explain why that potential builds up; only why that concentration ratio establishes: the Nernst law alone does not explain the effect. Furthermore, the statement is valid only for a well-defined semipermeable membrane. We can add that the same concentrations develop if initially identical concentrations are present on both sides. The finite width of the membrane creates its electric potential, which is a consequence of charge separation, as stated. If there are no ions initially inside, then no charge separation occurs, and therefore, no ion transport begins due to the membrane’s electric potential. However, the thermal driving force exists, and the started ion transport quickly builds the electric layer, which effectively helps to establish the resting potential.
This robustness also explains why, during replication (cell division), cells also inherit their operating ability and resting potential; the existing membrane cell continues to grow in the same way. Its voltage does not depend on its size, so it remains constant, and the voltage defined in this way adjusts the concentrations accordingly. Moreover, it explains why, during evolution, cells first formed: a couple of lipids came together to form an imperfect membrane with holes, establishing cellular electricity. As Fig. 9 depicts, the concentrations may adapt to the living conditions (whether living in seawater), but the general operating principles remain the same. The seawater environment enables the use of thinner membranes and lower membrane voltages; however, the same electric field is used in all cases.
Given that neurons share the external (global) concentrations (defined by the vast amount of ions in bulk), they provide a firmly fixed offset value to the always-the-same electric potential. This way, the operations of the individual neurons do not interfere. Furthermore, the internal (local) concentrations can adjust themselves even following a very rough charge perturbation while issuing an AP; moreover, when the living organism is growing.
5.3 Goldman-Hodgkin-Katz potential
It was already stated, based on experimental evidence, that "the membrane permeability to the ions has nothing to do with the potential generation and the ions adsorption on the membrane surface generates the membrane potential", for a review see [2]. The idea itself is nonsense: in a balanced state, no ion transport happens, so even without permeability, the balanced state persists; the resting potential has nothing to do with either ions’ mobility or permeability nor with ion absorption [20]. Furthermore, there is no idea in GHK about whether the ’setpoint’ (why that specific concentration or potential difference) is present and why the same potential is reset after rough perturbations such as issuing an AP or replicating a cell. The causality is reversed: the potential is a static concept and is set electrically, and the two gradients form the experienced concentrations from the available solvent molecules. For the time course of forming the gradients, mobility and permeability play a role, but not in determining the concentrations or the resting potential.
In our clear physical picture, the thermodynamic forces on the one side and on the other side of the membrane are summed. They counterbalance each other; furthermore, jointly the effect of the neuronal condenser, see Fig. 9. As we emphasized, the concentration gradients are ion-specific. Furthermore, as we discussed in connection with Eq.(2), the Nernst equation comprises a per-ion indefinite constant (a potential difference). To calculate a linear combination of terms comprising an arbitrary constant is nonsense, and so is adding absolute concentrations on the different sides of the membrane, or changing the base of the logarithm used in a calculation to match the experimental value. It is not more than number magic. The potential is described by coupled equations as discussed in connection with Eq. (39).
The $Ca^{2+}$ ions do not fit into the GHK picture. One of the reasons why GHK cannot be good is that $Ca^{2+}]$ is omitted. Biologically, it is hard to believe that $Ca^{2+}$ does not participate in the game of life (especially since biology sees the need for $Ca^{2+}$ pumps). Physically, one concentration alone on one side of the membrane cannot keep balance; as the different ion exchange processes, Table 1 and Fig. 9 show. When adding a new ion to the solution, the sum concentration increases, and so the electrical force increases, forcing the previous elements to find new concentrations on both sides. The appearance of a new chemical element indirectly changes the concentrations of the others (a good example is the role of the negligible amount of $Ca^{2+}$).
@@ 6 Summary
We provided a holistic, multi-level view of neuronal operation and described quantitatively the actual physical processes (instead of biophysics’ fairy tales about hypothesized lipid mechanisms); this way, we provided the promised ’new understanding’ (and the need for rewriting books in neuroscience). The classical neuron model, despite its ambition to be an electrical model, contradicts fundamental laws of electricity and conservation laws of science; furthermore, it cannot interpret other (thermodynamic, mechanical, optical) disciplinary phenomena. The thermodynamic model(s), on the other hand, cannot explain the electrical features of the observed processes. We proved that no compromise is needed: the pressure and the voltage observe the same physical process by different means. Neuroscience forgot that the neural current consists of ions, and they repel each other, in a confined volume.
In simple words, the neuron combines different disciplines; it works like a two-tact biological internal combustion engine, and can be described as a special Carnot-type thermodynamic engine. The synaptic input currents operate the ignition. The influx of $Na^{+}$ ions acts like an explosion, producing a vast pressure impulse, due to the electric repulsion of ions. The neuron’s volume remains practically unchanged, and the pressure wave moves the electrolyte (containing dissociated ions) out of the neuron through the only available output channel: the axon. In the first stroke, the pressure impulse invests energy into compressing the elastic membrane and starting a compression wave (soliton). In the second stroke, the elastic membrane pumps out electrolyte through the AIS, which "resists", and this way generates a potential wave, known as AP. The compression produces heat, observable as a temperature increase, and the expansion consumes heat, observable as a temperature decrease; these are well-understood processes in thermodynamics. The elastic membrane functions as a damped oscillator, which, in its negative amplitude phase, reverses the direction of electrolyte flow (and, consequently, the direction of the current carried by the ions). In the discipline of electricity (when using the serial $RC$ model), it can be interpreted as a capacitive current, which in physiology is observed as hyperpolarization that inverts the polarity of the output voltage. The ions represent both mass and charge simultaneously, and correspondingly, in addition to the electric potential wave, mechanical, optical, and other feature changes can be observed. In different disciplines, pressure and voltage represent the same phenomenon, using concepts from various disciplines.
Research already provided the fundamental principles of a correct new ‘net electric’ neuron model, the laws of motion for describing biology, and provided initial steps toward a correct cross-disciplinary model (promising a unified and scientifically complete electric/thermodynamic model); all these open new perspectives for making exploratory experimental and theoretical research. The perspective includes an understanding of the actual thermodynamic and mechanical behavior of cells (and possibly, life) and the relationship between neural information and its carriers. The fact that the neuron receives through its synapses less than 3% of the energy it consumes during its AP cycle underscores that a neuron is a “summing amplifier” rather than a “repeater”. The time window is very narrow compared to the length of the operating cycle, so the neuron receives inputs only from a small subset of neurons connected to its synapses (the late input signals are gated, as Fig. 12 shows).
@@
We have shown that the neuron represents a well-designed, simple control circuit with separate ground and excited state mechanisms and established that the neuron works as a simple control circuit where the construction defines the setpoint and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. The geometry and the overall concentration of the cellular fluids define the nominal values of the electric and thermal gradients. Our results are in line with the conclusions "It is shown theoretically that the membrane potential enhancement becomes greater with decreasing the electrolyte concentration and pore radius" [19]. The ion currents deliver ions through the permanently open "resting" and the gated "transient" ion channels, respectively, which means they affect the gradients until the equilibrium state is kept or restored. By understanding that the biological "construction" (the thickness of the isolating semipermeable membrane and the presence of ion channels in the walls of the membrane separating ionic solutions) defines the value of the resting potential, we defied the decades-old fallacy that some mystic concerted operation of ion channels and some linear combination of purely-defined permeability values (the GHK equation) define the resting potential of neurons. By using the principles of classic electricity, we calculated, that neurons handle charge as electricity describes (electrolyte) condensers. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat absorption" by neurons and the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH claimed in the physical picture behind their empirical description and mathematical equations, and introduced the correct physical model by providing its detailed mathematical description.

### v3

@@ 7.2 Heat emission/absorption
In the framework of the classic theory, Ohmic currents flow through resistors that dissipate heat due to friction, no matter in which direction ($W=I^{2}\times R$) the ion currents flow. The existence of "cooling", alone, undermines the credibility of the classic theory [66]. The issue here is that the dissipation was attributed to the "leakage current." As Hodgkin wrote: "Hill and his colleagues found [22] that an initial phase of heat liberation [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." [83] "All authors came to similar conclusions: during the action potential, no significant net heat is produced. Transient heat releases are mostly reabsorbed in the second phase of the action potential. …The finding of a reversible heat release during the action potential of nerves is a striking and very fundamental fact. It is inconsistent with the HH model. The physics underlying the nervous impulse must rather be based on reversible processes". [84]
Heat liberation and absorption are direct experimental evidence of the "slow current" and the mechanisms described above. The lack of leakage current (see section 4.5.4) not only validates our model, but also solves the long-standing mystery of reversible heat release during the AP. Pressure changes accompany the electrical process and naturally account for the observed temperature changes. Our Speculations in Appendix B on the energy consumption of the nervous impulse underpins that the complex physical process comprises mostly reversible disciplinary processes: most of the energy of AP is stored as reversible elastic potential energy. Hence, its dissipation is also negligible: "no significant net heat is produced" [84]. (Alternatively, it may be hypothesized that the positive $K^{+}$ ions, while traveling ahead of the positive $Na^{+}$ ions under the effect of a potential-free space in the second phase of AP, not only turn with their positive charge the offset voltage to negative, but on the way the $K^{+}$ ions collect the heat dissipated by the $Na^{+}$ ions. "Laws of science cannot describe life".)
The shocking conclusion in [63], that "brain cells communicate with mechanical pulses, not electrical signals", more precisely sounds that mechanical pulses (pressure waves) deliver the charge carriers of the electrical signals, in close interaction with the repulsion of the delivered charged carriers. In Equation (7), we also must consider the mechanical force delivered by the pressure wave. Furthermore, the nervous pulse transmission over the axons must be revisited. The classical mechanism based on coordinated action of ion channels is undoubtedly wrong. The enormous pressure, the incompressible electrical fluid, and the repulsion of the delivered particles suggest an alternative and more reasonable mechanism.
@@ 7.3 Neural entropy
The theoretical diagram line in Fig. 15, based on the presented correct physical model, might serve as a good starting point for interpreting entropy for neuronal operation, again in analogy with the Carnot cycle; although the charge of the ions (the macroscopic current) complicates the process. Furthermore, material convection happens in different ill-defined volume parts of the membrane in different directions. Initially, in the resting state, the ions are in a (relatively) disordered state. After the rush-in, in the dynamic layer, the ions have a highly ordered macroscopic speed (see Stokes-Einstein speed, the gradually changing gradients, and the finite speed) component toward the AIS. In the layer, the maximum ordered state is reached when the voltage drop on the AIS reaches its peak value (the current is slow). Then, the ions slow down, and at zero voltage gradient, they reach again a minimally ordered state. After that, due to the ’capacitive current’, the current reverses its direction, and the backward current again represents an ordered state. Perhaps the latter-mentioned negative contribution is the one referred to as ‘negative entropy’ when describing life in terms of thermodynamic processes? We must be aware that, in the background, energy-producing processes are at work, involving the ordered transport of material. Considering the neuron, based only on the described primary processes, would be a mistake.
@@
The above discussion might help in understanding the physical entropy in the neuron; although, as discussed above, there are issues with calculating partial derivatives (and so: deriving entropy from the Gibbs energy of the system) and with the applicability of Boltzmann assumptions to ensembles of ions. However, the information entropy is different. As our discussion underpins, the "information" arrives at the neuronal inputs as charge carriers, and the time course of the input currents, in cooperation with the neural environment and the neuron itself, carries the information in some form. The information is distributed in space and time [85, 86], so we can be sure that using the pulse count as a measure of neural information is wrong. We must distinguish the carrier of the information, the ion, and the information content that its temporal appearance carries. After having the correct physical model of operation, we have a chance to gain more knowledge about how the information is encoded in the spatially and temporally distributed observable signals; furthermore, we have greater hope of understanding neural information processing, including the brain’s entropy handling.
@@ 8 Summary
The paper reviewed the competing disciplinary electrical and thermodynamic theories for describing neuronal operation and described the non-disciplinary unified theory that resolves all former controversies between the two. We have shown that the recent incompatibility between the electrical and thermodynamic theories is the restrictive disciplinary approach on both sides, and that using a cross-disciplinary approach (and revisiting the fundamental laws of science for describing living matter) perfectly describe neuronal operation. We have shown that the interplay of those two disciplines is vital in all aspects of neuronal operation. The neuron represents a well-designed, simple control circuit with separate ground and excited state mechanisms, where the construction defines the setpoint and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. The geometry and the overall concentration of the cellular fluids define the nominal values of the electrical and thermal gradients. The AP is part of the operation of the control circuit. The low and largely variable speed of ionic current, in which the discrete features of the carrier dominate, creates local ionic gradients and movement in the electrolyte in the segments of the neuron.
@@
Slow currents deliver ions through the permanently open "resting" and the gated "transient" ion channels, respectively, which means they affect the gradients until the equilibrium state is kept or restored. By introducing the theoretical idea that biological currents are slow, we underpinned the experimental observation by Hodgkin and Huxley [74] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". By using the principles of classical electricity, we calculated, that neurons handle charge as electricity describes (electrolyte) condensers. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat production" by neurons and the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH claimed in the physical picture behind their empirical description and mathematical equations, and introduced the correct physical model by providing its detailed mathematical description. By understanding that the biological "construction" defines the value of the resting potential, we defied the decades-old fallacy that a mystic concerted operation of ion channels and a linear combination of poorly-defined permeability values (the GHK equation) define the resting potential.